When a thin film of a transparent material of refractive index 1.5 for wavelength 5890 Å is inserted in one of the arms of a Michelson's interferometer, a shift of 65 circular fringes is observed. Calculate the thickness of the film.
1
Expert's answer
2020-01-06T10:15:23-0500

For Michelson interferometer without the film

"\u2206_1=2dn_0\u20142tn_0=m_1\u03bb (1)"

with the film

"\u2206_2=2ln+2(d-l)n_0-2tn_0=m_2\u03bb (2)"

where d and t is Michelson interferometer arms lengths


Using (1) and (2) we got

"\u2206=\u2206_2-\u2206_1=2ln-2ln_0=k\u03bb (3)"

In our, case we have k=65, λ=5890×10-10 m, n=1.5, n=1

l=38.2×10-6 m


Answer

38.2×10-6 m